in joules experiment when the current is doubled the temperature of the liquid will be

Grade:12

1 Answers

Subhadeep pan
5 years ago
The relation is:-
Current \alpha  1/\sqrt{}T    (if p is considered to be constant)
So if current is doubled then temperature will be 2/\sqrt{}T